OpenStudy (anonymous):

Shane has a segment with endpoints C(3, 4) and D(11, 3) that is divided by a point E such that CE and DE form a 3:5 ratio. He knows that the distance between the x-coordinates is 8 units. Which of the following fractions will let him find the x-coordinate for point E? 3 over 5 5 over 3 3 over 8 5 over 8

OpenStudy (anonymous):

@Er.Mohd.AMIR Thank you

OpenStudy (amistre64):

thats incorrect, work thru the method and youll see why

OpenStudy (amistre64):

if you are just going to accept any direct answer given to you, then you will find that you receive wrong answers more times than not.

OpenStudy (anonymous):

@amistre64 Could you help me through it then?

OpenStudy (amistre64):

sure tell me what you know or think about it first, lets see what ideas you have to work with.

OpenStudy (anonymous):

I think that the answer could be 3/8 but i am very confused

OpenStudy (amistre64):

you are right, but the question itself is not well stated. what is your reasoning for 3/8 ?

OpenStudy (anonymous):

Because of the 3:5 ratio it would add it up too 8 total and E would fall at three. Does that make sense?

OpenStudy (amistre64):

that makes perfect sense. but i think your E=3 is off let me dbl chk we want to add 3 to the C x component; E x component is 3+3

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

now the thing is bad becuase, we can use 5/8 as well 8/8 = 1 5*1 = 5 subtracting from the right side: 11-5 = 6 is just as valid

OpenStudy (amistre64):

so in my view, the question has 2 right answers.

OpenStudy (anonymous):

Oh so its kind of a gamble then?

OpenStudy (amistre64):

its a bad question if you can only pick one :) they most likely want 3/8

OpenStudy (anonymous):

Thank you so much I really appreciate it

OpenStudy (amistre64):

youre welcome and good luck. if you can give feedback to your teacher, advice them of the problem and they can make it better with any luck

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

@mathisfun007 What?